1.3 The following table lists foreign exchange rates between U.S. dollars and British pounds during April Date U.S. Dollars per GBP 1.9564 1.9293 1.914 1.9374 1.961 1.8925 1.8822 1.8558 1.796 1.7902 1.7785 Date 4/18 4/19 4/20 4/21 4/22 4/25 4/26 4/27 4/28 4/29 U.S. Dollars per GBP 1.7504 1.7255 1.6914 1.672 1.6684 1.6674 1.6857 1.6925 1.7201 1.7512 4/4 4/5 4/6 4/8 4/12 4/13 4/14 4/15 Which day would have been the best day to convert $200 into British pounds? Which day would have been the worst day? What would be the difference in pounds?
Answer #1

best day would be 4/25 as the no. of pounds will be = 1*200/1.6674 = 119.9472

worst day would be 4/7 as the no. of pounds will be = 1*200/1.961 = 101.9888

The difference in pounds = 119.9472 - 101.9888 = 17.9584